# GenOM: Ontology Matching with Description Generation and Large Language Model

### v1

@@ 4.1 Datasets and Evaluation Metrics
The experiments were conducted on the OAEI 2024 Bio-ML [^4] https://krr-oxford.github.io/OAEI-Bio-ML/2024/index.html track [25] whose benchmarks are designed for biomedical ontology alignment tasks. The dataset comprises five sub-tasks involving six widely used biomedical ontologies: Systematized Nomenclature of Medicine - Clinical Terms (SNOMED-CT), National Cancer Institute Thesaurus (NCIT), Foundational Model of Anatomy (FMA), Human Disease Ontology (DOID), Orphanet Rare Disease Ontology (ORDO), and Online Mendelian Inheritance in Man (OMIM). The two official evaluation protocols of the Bio-ML track were adopted: global matching which focuses on ranking the correct target concept among a list of candidates, and local ranking which is to evaluate the system’s ability to identify correct mappings among all the possible concept pairs across two ontologies. Precision (P), Recall (R), and F1-score are measured for global matching, and Mean Reciprocal Rank (MRR) and Hit@1 are calculated for local ranking. These metrics provide a comprehensive view of the OM systems.
@@ 4.2 Experiment Setup
To systematically evaluate the proposed framework, we designed distinct experimental settings tailored to each module of the pipeline. Definition Generation: For semantic enrichment, the Qwen2.5-7B-Instruct-1M [^5] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct-1M LLM was utilised to generate concise definitions based on concept-level contextual information. The generation process was controlled with a temperature of 0.7 and a top_p value of 0.9, ensuring that the definitions remained focused and semantically aligned with the underlying concepts. Candidate Mapping Generation: Cosine similarity computations and HNSW-based indexing were implemented via the faiss library to efficiently retrieve the top-$10$ most similar concepts for each source entity. LLM-based Judgement: The same Qwen2.5-7B-Instruct-1M model was applied to perform binary equivalence classification over the candidate pairs. To accelerate inference, we adopted the float16 data type. Post-processing and Result Fusion. In the final stage, results were filtered using a token probability threshold $\lambda_{\text{prob}}$ of 0.99 and a cosine similarity threshold $\lambda_{\text{cs}}$ of 0.97. These thresholds were initially optimised on the SNOMED-NCIT (neoplas) task, then fixed and consistently applied across all the other tasks. Moreover, since BERTMapLt demonstrated stronger performance on the neoplas task among the exact matching models, it was selected for application to the remaining tasks as well.
@@ 4.3 Overall Results
Table 5: Overall performance on the five OAEI 2024 Bio-ML tasks. NA indicates no results as the systems do not supporting the calculation of the metrics. Bold indicates the best performance, while underline indicates the second-best. All the baseline results are from the track’s website, where the LogMapLt and the BERTMap series are reimplemented with system submissions, while the other baselines submitted the final result files without system reimplementation, probably using settings optimised for each task.
Table 5 presents the comparison between our proposed method and several state-of-the-art ontology alignment systems on the Bio-ML track. The results demonstrate that our model achieves consistently strong performance across all tasks, ranking among the top three in every case. Moreover, in tasks where our method ranks second, the performance gap with the best-performing system is marginal—for instance, only 0.006 in the SNOMED-NCIT (pharm) task and as small as 0.001 in the SNOMED-FMA (body) task.
To ensure a fair assessment and avoid overestimating the performance of our model, we excluded the neoplas task from the overall evaluation, as it was used during threshold tuning. When evaluated solely on the remaining unseen tasks—where no threshold optimisation was performed—our framework still achieved the highest average F1 score of 0.769. This surpasses the second-best method, BERTMap (0.762), and the third-best, LogMapBio (0.760). These results indicate that the proposed approach not only performs competitively on tuned datasets but also maintains strong and consistent performance across previously unseen tasks, highlighting its robustness and generalisability in biomedical ontology alignment.
In addition, when compared to the other LLM-based OM system LLM4OM which leverages ChatGPT-3.5 and OpenAI’s embedding model as reported in its original paper, our approach GenOM delivers consistently better performance across all evaluated tasks, including the last four tasks where GenOM generalises the hyper parameter settings optimised from the first task.
4.4 Ablation Study
4.4.1 Impact of Definition Enrichment on Local Ranking and Candidate Retrieval

### v2

@@ 4.1 Datasets and Evaluation Metrics
The experiments were conducted on the Bio-ML [^6] https://krr-oxford.github.io/OAEI-Bio-ML/2024/index.html track [37] of the Ontology Alignment Evaluation Initiative (OAEI) in 2024, which provides benchmarks specifically designed for biomedical ontology alignment, with a particular focus on machine learning-based systems. This track comprises five sub-tasks involving six widely used biomedical ontologies: Systematized Nomenclature of Medicine Clinical Terms (SNOMED-CT), National Cancer Institute Thesaurus (NCIT), Foundational Model of Anatomy (FMA), Human Disease Ontology (DOID), Orphanet Rare Disease Ontology (ORDO), and Online Mendelian Inheritance in Man (OMIM).
The two official evaluation protocols of the Bio-ML track were adopted: global matching which is to evaluate the system’s ability to identify correct mappings among all the possible concept pairs across two ontologies, and local ranking which focuses on ranking the correct target concept of a given source concept among a list of candidates.
Precision (P), Recall (R) and F1-score are measured for global matching, and Mean Reciprocal Rank (MRR) and Hit@1 are calculated for local ranking. In the competition, these metrics provide a comprehensive assessment of the OM systems.
All evaluations strictly follow the official OAEI Bio-ML evaluation protocols and metric definitions, using the provided reference alignments, to ensure fair and comparable results across the systems.
@@ 4.2 Experiment Setup
To systematically evaluate the proposed framework, we designed experimental settings tailored to each module of the pipeline.
Definition Generation: For semantic enrichment, Qwen2.5-7B-Instruct [^7] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct [^8] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct, Qwen2.5-32B-Instruct [^9] https://huggingface.co/Qwen/Qwen2.5-32B-Instruct and Llama3.1-8B-Instruct [^10] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct were utilised to generate concise definitions based on concept-level contextual information. For all definition generation experiments, we fixed the decoding hyperparameters to temperature = 0.7 and top_p = 0.9 across all the LLMs. This setting corresponded to a moderately conservative sampling regime, which reduced excessive stochasticity in generation while still allowing limited variability beyond fully deterministic decoding. Such a configuration is suitable for definition generation, where outputs were expected to be stable and informative rather than highly diverse.
It is important to note that these hyperparameters were kept fixed throughout the experiments rather than tuned for individual models or tasks. This design choice allowed us to isolate the effect of the LLM itself on definition quality, avoided coupling the results to model-specific decoding behaviour, and ensured comparability across different models and experimental settings.
Candidate Mapping Generation: Cosine similarity computation and HNSW-based indexing were implemented via the faiss library to efficiently retrieve the top-$10$ most similar concepts for each source entity.
LLM-based Equivalence Judgement: The Qwen2.5-32B-Instruct LLM was applied to perform binary equivalence classification over the candidate pairs.
Post-processing and Result Fusion: In the final stage, we applied two decision thresholds: the cosine similarity threshold ($\lambda_{\text{cs}}$) and the LLM-based probability threshold ($\lambda_{\text{prob}}$). To determine suitable values, we considered a predefined set of candidate thresholds $\{0.5,0.6,0.7,0.8,0.9\}$ with a step size of $0.1$. This range was chosen to focus on practically meaningful operating points and to exclude extreme high threshold values, which tend to produce overly conservative decisions and poor cross-task robustness in ontology matching. Threshold selection is performed on a single reference task, SNOMED-NCIT (Neoplas). Based on its development results, a unified threshold value of $0.9$ was selected for both $\lambda_{\text{cs}}$ and $\lambda_{\text{prob}}$. These values were then fixed and applied unchanged to all remaining ontology matching tasks. Under such a setting, GenOM may not achieve the optimum performance on other tasks beyond SNOMED-NCIT (Neoplas), but we believe this setting is more consistent with real-world deployment and can reflect the framework’s real performance especially on its robustness on different tasks. In addition, results based on optimal thresholds obtained through task by task tuning are reported in Appendix A, for post-hoc analysis as a reference of GenOM’s optimum performance and for comparison with other systems’ results produced under this optimal setting.
With regard to the use of definition information, various LLMs were employed to generate textual definitions for medical concepts. However, in order to maintain experimental consistency, all components of the study that make use of definitions, except for the section specifically evaluating the definition quality produced by different LLMs, use a unified set of definitions generated by the Qwen2.5-7B-Instruct model.
For ease of reference, we use the following abbreviations: Qwen7B (Qwen2.5-7B-Instruct), Qwen14B (Qwen2.5-14B-Instruct), Qwen32B (Qwen2.5-32B-Instruct), and Llama8B (Llama3.1-8B-Instruct). Additionally, we write GenOM (Qwen32B) for the GenOM framework using Qwen32B as the LLM-based alignment model, i.e., the component responsible for predicting equivalence correspondences.
@@ 4.3 Overall Results
Table 5: Overall performance on the five OAEI 2024 Bio-ML tasks. NA indicates no results as the systems do not support the calculation of the metrics. Bold indicates the best performance, while underling indicates the second-best performance. All the baseline results are from the track’s website, where the LogMapLt and the BERTMap series were reimplemented with system submissions, while the other baselines submitted the final result files without system reimplementation, probably using settings optimised for each task. Results of LLM4OM, LogMap-LLM, and Hu et al. were taken from their original papers. LLM4OM and Hu et al. (2025) used fixed thresholds across tasks, while LogMap-LLM relied on OpenAI and Gemini models without an explicit threshold parameter.
Table 5 presented the performance of our method on the Bio-ML track in comparison with several state-of-the-art ontology alignment systems. These systems can broadly be grouped into three categories: (1) traditional ontology matching methods, LogMap [8], LogMapBio [38], LogMapLt [38], and Matcha; (2) pre-trained language model (PLM)-based approaches, BERTMap [15], BERTMapLt, and BioSTransMatch [23]; and (3) LLM-based methods, Hu et al. (2025) [27] (LLaMA-3.3-70B-Instruct), LLM4OM [10] (ChatGPT-3.5), and LogMap-LLM [28] (GPT-4o).
Overall, GenOM (Qwen32B) exhibited competitive performance relative to other baseline systems on the Bio-ML track.
  • On the SNOMED–NCIT (Neoplas) task, GenOM (Qwen32B) achieved the best results across all three evaluation metrics, namely F1-score, MRR, and H@1. In particular, its F1-score exceeded that of LogMapBio, a LogMap variant that explicitly integrated additional biomedical knowledge.
  • For the SNOMED–NCIT (Pharm) task, GenOM (Qwen32B) attained the highest F1-score among all evaluated systems. Although GenOM ranked second with respect to MRR and H@1, the absolute differences to the top-performing method, BERTMap, were small, amounting to 0.004 for MRR and 0.008 for H@1.
  • On the SNOMED–FMA (Body) task, GenOM (Qwen32B) ranked second in terms of both MRR and H@1, with marginal differences compared to the top-ranked system. While GenOM achieved the third-highest F1-score, the absolute gaps to the first- and second-ranked methods were limited. Specifically, BERTMap attained the highest F1-score of 0.790, whereas GenOM achieved 0.771. By contrast, the fourth-ranked system, LogMapLt, recorded an F1-score of 0.696. This corresponds to an absolute F1-score difference of 0.075 compared to LogMapLt, indicating that GenOM achieved performance comparable to the leading methods on this task.
  • On the OMIM–ORDO task, GenOM (Qwen32B) achieved the best performance across all three evaluation metrics. In particular, it outperformed the second-best system on ranking-based metrics, exceeding BERTMap by 0.05 in MRR and by 0.074 in H@1.
  • For the NCIT–DOID task, GenOM (Qwen32B) again achieved the highest scores on MRR and H@1. Although it ranked second in terms of F1-score, the absolute difference to the top-ranked system was only 0.007, further indicating that GenOM remained competitive on this task.
Compared with existing LLM-based approaches, GenOM (Qwen32B) achieved strong overall performance while relying on a model with a smaller parameter size. On the two Bio-ML tasks where Hu et al. (2025) [27] reported results (SNOMED–FMA and OMIM–ORDO), GenOM (Qwen32B) increased the F1-score by 0.518 and 0.107, respectively (averaging a 0.313 absolute gain across the two tasks). Across all five Bio-ML tasks, GenOM (Qwen32B) improved the mean F1-score over LLM4OM by 0.205 (35.3%) and over LogMap-LLM by 0.071 (9.88%).
In general, GenOM (Qwen32B) exhibited consistently strong and stable performance in different tasks and evaluation metrics. Even in tasks where it did not achieve the top rank, the performance gaps of the best-performing systems remained small, indicating that GenOM (Qwen32B) operated at a level comparable to state-of-the-art approaches. Furthermore, its strong performance on local ranking metrics demonstrated that GenOM (Qwen32B) was effective in distinguishing semantically similar biomedical concepts, highlighting its effectiveness in fine-grained semantic discrimination.
5 Definition Generation Analysis
Concept LLM-generated definition (label-like)
Middle cerebral artery infarction Infarction of the middle cerebral artery.
Immune system organ benign neoplasm A benign neoplasm originating in an organ of the immune system.
Table 6: Examples of label-like definitions that largely restate the concept label without providing additional semantic information.
In this section, we analyse the quality of the definitions generated by different LLMs during the Definition Generation stage. Several LLMs were used to produce definitions for the same set of biomedical ontology concepts, enabling a comparative study of their ability to capture domain semantics and support downstream ontology alignment.
Accurate and informative definitions play an essential role in ontology matching, particularly in the biomedical domain where many concepts differ only subtly [26]. The generated definitions in GenOM are expected to be factually accurate and medically informative, while being sufficiently discriminative to differentiate the target concept from closely related alternatives.
In our experiments, we observed that definitions generated by different LLMs vary substantially in quality, and that simple surface properties such as length or lexical overlap provide limited insight into their quality. These observations motivate the need for an evaluation framework that can assess both the correctness of the generated content and its ability to discriminate between semantically similar concepts, while avoiding artificially high scores for trivial or label-like rewritings as illustrated in Table 6.
5.1 Hybrid Evaluation Method
Algorithm 1 Hybrid Definition Quality Scoring Method with Nearest-Neighbour Contrast
1:Concept label $A$, generated definitions $Def=\{def_{1},\dots,def_{n}\}$, ontology $\mathcal{O}$ containing AA, SapBERT encoder $f(\cdot)$, penalty coefficient $\lambda$ (set to 11 in our experiments).
2:Final adjusted scores $S=\{s_{1}^{\prime},\dots,s_{n}^{\prime}\}$.
3:Nearest-Neighbour Selection (Contrastive Pair Construction):
4:  Encode the label of AA using SapBERT: $\mathbf{e}_{A}\leftarrow f(A)$.
5:  For each candidate concept $x\in\mathcal{O}\setminus\{A\}$, compute $\mathbf{e}_{x}\leftarrow f(x)$ and similarity
6:  $\mathrm{sim}(\mathbf{e}_{A},\mathbf{e}_{x})$.
7:  Select the nearest neighbour $B\leftarrow\arg\max_{x\in\mathcal{O}\setminus\{A\}}\mathrm{sim}(\mathbf{e}_{A},\mathbf{e}_{x})$.
8:  Note: AA and $B$ are retrieved from the same ontology $\mathcal{O}$, and the retrieval uses
9:  only lexical information (e.g., labels), not generated definitions.
10:LLM-based Scoring: Query GPT-5-mini once, providing label $A$, nearest neighbour $B$, and the set of definitions $\{def_{1},\dots,def_{n}\}$ in a single prompt (see Appendix B.2).
11:  The LLM returns, for each $def_{i}$, a pair of scores:
12:   Discriminative score $d_{i}\in\{1,\dots,5\}$ (how well $def_{i}$ distinguishes $A$ from $B$).
13:   Correctness score $c_{i}\in\{1,\dots,5\}$ (semantic correctness of $def_{i}$).
14:for $i=1$ to $n$ do
15:  Normalised LLM Score: $s^{\text{norm}}_{i}\leftarrow\frac{d_{i}+c_{i}}{10}$.
16:  Jaccard-based Penalty:
17:  Tokenise label AA into $T_{A}$, definition $def_{i}$ into $T_{i}$. Compute $J_{i}\leftarrow\frac{|T_{A}\cap T_{i}|}{|T_{A}\cup T_{i}|}$.
18:  if $J_{i}>0.5$ then
19:   $P_{i}\leftarrow\lambda\cdot(J_{i}-0.5)$
20:  else
21:   $P_{i}\leftarrow 0$
22:  end if
23:  Final Score: $s_{i}^{\prime}\leftarrow s^{\text{norm}}_{i}-P_{i}$.
24:end for
25:return $S=\{s_{1}^{\prime},\dots,s_{n}^{\prime}\}$
We employ GPT-5-mini [^11] https://platform.openai.com/docs/models/gpt-5-mini, a proprietary LLM with a larger parameter scale than the open-source LLMs used in GenOM, as an external evaluator for assessing the quality of generated definitions. The evaluator was not used in any stage of definition generation or ontology matching, thereby serving solely as an independent assessment component.
The proposed evaluation framework consists of three components: (i) Nearest-Neighbour Selection for constructing contrastive concept pairs, (ii) LLM-based Scoring of correctness and discriminativeness, and (iii) a Jaccard-based Penalty to reduce redundancy with the concept label. The complete procedure is summarised in Algorithm 1.
Concretely, for each concept $A$ and its set of generated definitions $\{def_{1},\dots,def_{n}\}$, we first identify a lexically nearest neighbour $B$ from the same ontology using SapBERT [39] embeddings. This neighbour serves as a contrastive reference for evaluating the discriminative quality of definitions. The retrieval relies solely on lexical information (e.g., concept labels), and the nearest neighbour is selected as the top-1 most similar concept within the same ontology. Importantly, generated definitions are not used during this retrieval step, ensuring that the selection of $B$ is independent of generated definitions.
Next, the concept label $A$, its nearest neighbour $B$, and all candidate definitions are provided to GPT-5-mini in a single prompt (see Appendix B.2). For each definition $def_{i}$, the LLM assigns two scores on a discrete scale from 1 to 5: a correctness score $c_{i}$, reflecting semantic and biomedical validity, and a discriminativeness score $d_{i}$, measuring how effectively $def_{i}$ distinguishes $A$ from $B$. These scores are combined into a normalised LLM score, as defined in Algorithm 1, yielding values in the range $[0,1]$ as a relative quality score.
To penalise definitions that closely resemble the concept label, we compute a Jaccard similarity $J_{i}$ between the tokenised label $A$ and the definition $def_{i}$. Both the label and the definition are tokenised at the word level, and common English stopwords (e.g., a, an, the, of, in) are removed prior to computing the intersection and union. This design ensures that the redundancy penalty is driven by semantically meaningful lexical overlap rather than superficial matches on high-frequency function words. When $J_{i}>0.5$, a penalty $P_{i}=\lambda(J_{i}-0.5)$ is applied (with $\lambda=1$ in our experiments); otherwise, the penalty is set to zero. The threshold of 0.5 was chosen as a conservative heuristic to penalise substantial lexical reuse, while allowing partial overlap that is often unavoidable in concise definitions. The final adjusted score is computed as $s_{i}^{\prime}=s_{i}^{\text{norm}}-P_{i}$, where higher values indicate higher-quality definitions.
Due to the computational cost of GPT-5-mini, we randomly sampled 100 concepts from each ontology for evaluation. For each concept, the number of generated definitions was fixed to four in our experiments, corresponding to definitions produced by different LLMs. The resulting score $S$ provides a composite measure that integrates semantic correctness, discriminative capability, and redundancy penalisation.
Beyond the automated evaluation method described above, we also conducted a small-scale manual inspection as a qualitative sanity check to provide additional context for the proposed evaluation method. Details of this analysis and its limitations are reported in Appendix B.1.
5.2 Results and Analysis for the Definition Evaluation Task
Table 7: LLM definition–generation performance across ontology subsets. Each subset contained 100 randomly sampled concepts. For each concept, the four LLMs independently generated a definition, and raw/final scores were averaged over the 100 samples. Note. Raw scores were obtained from the LLM-based assessment, and final scores incorporated the Jaccard-based redundancy penalty. Both scores were averaged over the 100 concepts per (task, ontology, LLM) combination.
Task Ontology LLM Avg. raw score Avg. final score
(0–1) (0–1)
Pharm SNOMED Qwen7B 0.768 0.752
Llama8B 0.744 0.721
Qwen14B 0.810 0.797
Qwen32B 0.801 0.794
NCIT Qwen7B 0.713 0.713
Llama8B 0.663 0.654
Qwen14B 0.765 0.764
Qwen32B 0.746 0.745
Neoplas SNOMED Qwen7B 0.776 0.775
Llama8B 0.745 0.707
Qwen14B 0.806 0.805
Qwen32B 0.813 0.809
NCIT Qwen7B 0.725 0.725
Llama8B 0.685 0.675
Qwen14B 0.785 0.784
Qwen32B 0.753 0.753
Body SNOMED Qwen7B 0.695 0.695
Llama8B 0.652 0.609
Qwen14B 0.676 0.674
Qwen32B 0.712 0.708
FMA Qwen7B 0.601 0.601
Llama8B 0.641 0.591
Qwen14B 0.724 0.695
Qwen32B 0.740 0.715
Disease NCIT Qwen7B 0.781 0.781
Llama8B 0.678 0.658
Qwen14B 0.752 0.750
Qwen32B 0.762 0.761
DOID Qwen7B 0.754 0.754
Llama8B 0.634 0.626
Qwen14B 0.755 0.750
Qwen32B 0.805 0.804
Disease OMIM Qwen7B 0.743 0.742
Llama8B 0.668 0.659
Qwen14B 0.740 0.735
Qwen32B 0.749 0.746
ORDO Qwen7B 0.802 0.802
Llama8B 0.725 0.721
Qwen14B 0.806 0.806
Qwen32B 0.804 0.803
Table 7 reported the average raw (LLM-based) and final (penalised) definition scores across ontology subsets, where the final score incorporated the Jaccard-based redundancy penalty. Three observations emerged.
First, the Qwen family consistently achieved higher final scores than Llama8B across the evaluated subsets. When averaging over all (task, ontology) subsets, Qwen32B and Qwen14B obtained higher mean final scores (0.764 and 0.756, respectively) than Qwen7B (0.734), while Llama8B yielded a substantially lower mean final score (0.662). Moreover, Qwen32B exhibited the strongest overall performance, achieving the best final score in five out of the ten subsets, with Qwen14B ranking first in four subsets. The only subset where a smaller model led was Disease–NCIT, where Qwen7B attained the highest final score (0.781).
Second, model scaling improved definition quality in general, but the gains were not strictly monotonic across subsets. Although Qwen32B achieved the strongest overall performance under the proposed evaluation, Qwen14B outperformed it on several individual subsets, including Pharm–SNOMED (0.797 vs. 0.794), Pharm–NCIT (0.764 vs. 0.745), Disease–ORDO (0.806 vs. 0.803), and Neoplas–NCIT (0.784 vs. 0.753). This showed that increasing model size did not consistently result in higher definition scores for every ontology subset.
[…]
Overall, these results suggested that larger Qwen models tended to produce higher-quality definitions under the proposed hybrid evaluation, with Qwen32B providing the most stable performance across subsets. At the same time, the performance gap between Qwen14B and Qwen32B remains small across most subsets, despite the general advantage of larger models. This pattern indicates that increasing model scale yields diminishing returns in definition quality, and suggest that model size, while influential, is not the sole factor determining final performance under this evaluation. The substantial difference between Llama8B and Qwen7B at a comparable parameter scale is consistent with this observation.
6 Impact of different LLMs
Figure 3: F1-Scores of GenOM under the LLM-only setting when the thresholds of cosine similarity and token probability are set to different values. Results of two matching tasks using Qwen7B and Qwen32B are reported.
Figure 3 presents F1-score heatmaps for two representative tasks, NCIT–DOID (Disease) and SNOMED–NCIT (Neoplas), evaluated across a grid of the token-probability threshold and the cosine-similarity threshold for both Qwen7B and Qwen32B. These results were based on the LLM-only setting, where the final alignment decisions were produced solely by the LLM without incorporating any exact-matching modules. Each candidate pair was evaluated using the labels, synonyms and generated definitions of the concepts, allowing us to directly compare the performance of different LLMs.
Across both tasks, Qwen32B consistently achieved higher F1 scores than Qwen7B, which is consistent with its larger model capacity. More importantly, the two models differ in their sensitivity to the cosine-similarity threshold.
For the NCIT–DOID task, both models showed very limited variation in F1 scores as the threshold changed across the examined range. This suggested that the alignment signals—derived from labels, synonyms, and generated definitions—were sufficiently strong and consistent that the equivalence-judgement stage of the LLM was largely insensitive to threshold selection. The underlying ontologies in this pair shared relatively homogeneous, clinically oriented concept representations, which made the equivalence classification problem comparatively straightforward even for the smaller model.
In contrast, the SNOMED-NCIT (Neoplas) task displayed substantial variability, particularly for Qwen7B.
This phenomenon was not confined to the SNOMED–NCIT (Neoplas) task; it also appeared consistently in other tasks involving SNOMED. When SNOMED participated in the alignment pair, the smaller model became highly dependent on strict cosine-similarity thresholds: relaxing the threshold resulted in a rapid increase in false-positive equivalence predictions, leading to sharply reduced F1 scores. This instability can be attributed to the difficulty of performing cross-ontology equivalence judgement when one side (SNOMED) encoded highly granular and structurally complex concepts.
SNOMED terms frequently contained fine-grained morphologic, anatomical or structural information, e.g. Adenolymphoma (morphologic abnormality), Dermatofibroma (disorder), Abducens nerve structure (body structure), whereas the corresponding concepts in NCIT were often expressed using more concise, high-level clinical terminology, e.g. Recurrent Respiratory Papillomatosis. Qwen7B struggled to reliably determine whether such asymmetric concept pairs should be treated as equivalent or merely related, producing many borderline cases that were filtered out only when strict cosine-similarity thresholds were applied.
By contrast, Qwen32B demonstrated far greater robustness across threshold settings. The larger model can better interpret SNOMED’s layered semantic structure and more reliably distinguish true equivalence from looser semantic relatedness, resulting in smoother heatmaps and reduced threshold sensitivity. These findings highlighted the intrinsic difficulty of aligning fine-grained, richly modelled ontologies such as SNOMED, and underscored the role of model capacity in achieving stable equivalence judgements under this setting.
7 Ablation Study
7.1 Impact of Definition Enrichment on Local Ranking and Candidate Retrieval